OpenStudy (anonymous):

Find the vertex and all intercepts of the parabola. Then sketch and determine window values. The problem is: f(x)=-x^2-2x+3

OpenStudy (anonymous):

the formula to find the mim/max point of a parabola is \[x=-\frac{b}{2a}\] so \[x = \frac{2}{-2} = -1\] so y = 4 so the vertex is (-1, 4) to find the intercepts just set x and y equal to 0 one by one the equation is y = -x^2 - 2x + 3 y = 0 0 = -x^2 - 2x + 3 x^2 + 2x -3 = 0 (x - 1)(x + 3) = 0 x = 1, -3 so the x intercepts are (1, 0) and (-3, 0) x = 0 y = 3 y intercept is (0, 3)